# Feynman Diagrams for Matter Wave Interferometry

### v2

@@ III.1.1 Diagram Values
where $T_{ij}$, $Q_{ijk}$, and $S_{ijkl}$ correspond to coefficients of the Taylor series expansion in space of the Earth’s gravitational field about the surface of the earth, where we assume a spherical earth with homogeneous mass density, and $z$ is the axis of the interferometer. More generally, these coefficients could correspond to the Taylor expansion of any spatially varying potential, such as that generated by a dipole beam, or nonuniform magnetic fields. With $L_{1}=0$, the interferometer phase can be computed by evaluating the 3D version of the spatial integrals of Sec II.2, which results in
@@ III.1.2 Validating with Numerics in 1D
To explore the extent to which these higher order quantum corrections under power law potentials associated with the spherical nature of the Earth could be measured with a state-of-the-art matter wave interferometer, we take the analytic expressions that come out of the diagrammatic formalism, and plug in typical operating parameters (parameters indicated in the caption of Table 1). In Table 1, we rank each of the terms that come out of this calculation in order of the size of their contribution to the interferometer phase shift, with the higher order quantum correction terms highlighted gray.
The largest term not captured by the semiclassical formalism has a relative magnitude of roughly one part in $10^{20}$ with respect to the largest overall $kgT^{2}$ term, which would be difficult to resolve experimentally with state-of-the-art atom interferometers. However, the numerical values of the Taylor series coefficients used above assume perfectly spherical earth with homogeneous mass density, and may differ significantly from realistic values. More realistic coefficients (accounting for the non-spherical nature of the earth, its inhomogeneous mass density, nearby structures, sea level variations, and other sources of gravitational potentials) could be orders of magnitude larger than those considered here [dubetsky_comment_2020, ufrecht_reply_2020]. We emphasize that the calculation presented here is intended as a demonstration of the method for computing quantum corrections to power-law potentials in general. In practice, the true coefficients ($T_{ij}$, $Q_{ijk}$, ...) would need to be measured at the site of the interferometer.
@@
It is important to carry out these calculations in three dimensions. Even though the semiclassical trajectories do not deflect in the directions transverse to the interferometer axis (i.e. the classical paths lie entirely along $z$), quantum corrections still depend on the transverse gravity gradients. For the potentials considered in Table 1, if the initial wavepacket is spherically symmetric ($w_{0,1}=w_{0,2}=w_{0,3}$), then the value of the 1-loop 3-point vertex diagram of Table 1 vanishes: the terms proportional to $Q_{xxz}$ (term 26) and $Q_{yyz}$ (term 27) cancel with the corresponding $Q_{zzz}$ contribution (term 23). This cancellation was noted in Ref. [ufrecht2020perturbative]. The diagrammatic formalism can be used to prove that, more generally, if the wavepacket is prepared with spherical symmetry, then the first-order quantum corrections to the semiclassical approximation vanish for any external potential $V[\mathbf{r}]$ that satisfies Laplace’s equation, $\nabla^{2}V[\mathbf{r}]=0$. We work out this result explicitly in Sec. E. Since the gravitational potential in mass-free regions satisfies Laplace’s equation, it follows that if the initial wavepacket is spherically symmetric, there are no quantum corrections to the interferometer phase shift to first order in $V[\mathbf{r}]$. For the numerical values of Table 1, we take the initial wavepacket to be twice as large along the interferometer axis than in the transverse dimensions, which gives rise to nonzero quantum corrections at first order in $Q_{ijk}$. The asymmetry in the initial wavepacket can be interpreted as arising from asymmetric cooling. For the parameters of Table 1, the momentum-space width is smaller along the interferometer axis than in the transverse directions, corresponding to a lower temperature along that axis. In Ref. [ufrecht2020perturbative], the authors compute the quantum corrections arising from a first-order expansion of the interferometer phase shift in $Q_{ijk}$. Using the same choice of initial wavepacket waists, $\beta_{1}=\beta_{2}=2\beta_{3}=\omega$, our first order result in $Q_{ijk}$ agrees with their first-order expression.
@@
Table 1: Phase shift terms for a Mach-Zehnder matter wave interferometer in 3D subject to gravity gradients associated with the spherical nature of the Earth, along with the diagrams which produce the term, ranked in order of the magnitude of their contribution to the total phase shift. The phase shifts are computed with the following parameters: $k_{z}=2\pi/(698\text{ nm})$, $T=1\text{ s}$, $g=9.8\text{ m/s}^{2}$, $T_{zz}=-2g/R$, $T_{xx}=T_{yy}=g/R$, $Q_{zzz}=6g/R^{2}$, $Q_{xxz}=Q_{yyz}=-3g/R^{2}$, $S_{zzzz}=-24g/R^{3}$, $S_{xxzz}=S_{yyzz}=12g/R^{3}$, $x_{0}=y_{0}=z_{0}=0$, $v_{z}=3\text{ m/s}$, $v_{x}=v_{y}=0$, $R=6.38\times 10^{6}\text{ m}$, $m=86.91\text{u}$, $v_{r}=\hbar k_{z}/m$, $\beta_{\ell}=2\hbar/(mw_{0,\ell}^{2})$ with $\ell=x,y,z$. $w_{0,z}=2\text{ mm}$, $w_{0,x}=w_{0,y}=1\text{ mm}$. The quantum correction terms are highlighted in light gray.
@@
2 $\frac{7}{12}\frac{m}{\hbar}gv_{r}T^{4}T_{\text{zz}}$ $1.58\times 10^{2}$ [Uncaptioned image]
3 $\frac{m}{\hbar}v_{r}v_{z}T^{3}T_{\text{zz}}$ $8.30\times 10^{1}$ [Uncaptioned image]
4 $\frac{1}{2}\frac{m}{\hbar}v_{r}^{2}T^{3}T_{\text{zz}}$ $9.10\times 10^{-2}$ [Uncaptioned image]
5 $\frac{31}{120}\frac{m}{\hbar}g^{2}v_{r}T^{6}Q_{\text{zzz}}$ $3.23\times 10^{-4}$ [Uncaptioned image]
6 $\frac{3}{4}\frac{m}{\hbar}gv_{r}v_{z}T^{5}Q_{\text{zzz}}$ $2.87\times 10^{-4}$ [Uncaptioned image]
7 $\frac{31}{360}\frac{m}{\hbar}gv_{r}T^{6}T_{\text{zz}}^{2}$ $7.17\times 10^{-5}$ [Uncaptioned image]
8 $\frac{7}{12}\frac{m}{\hbar}v_{r}v_{z}^{2}T^{4}Q_{\text{zzz}}$ $6.83\times 10^{-5}$ [Uncaptioned image]
9 $\frac{1}{4}\frac{m}{\hbar}v_{r}v_{z}T^{5}T_{\text{zz}}^{2}$ $6.38\times 10^{-5}$ [Uncaptioned image]
10 $\frac{3}{8}\frac{m}{\hbar}gv_{r}^{2}T^{5}Q_{\text{zzz}}$ $3.14\times 10^{-7}$ [Uncaptioned image]
11 $\frac{7}{12}\frac{m}{\hbar}v_{r}^{2}v_{z}T^{4}Q_{\text{zzz}}$ $1.50\times 10^{-7}$ [Uncaptioned image]
12 $\frac{1}{8}\frac{m}{\hbar}v_{r}^{2}T^{5}T_{\text{zz}}^{2}$ $6.99\times 10^{-8}$ [Uncaptioned image]
13 $\frac{3}{8}\frac{m}{\hbar}g^{2}v_{r}v_{z}T^{7}S_{\text{zzzz}}$ $8.81\times 10^{-10}$ [Uncaptioned image]
14 $\frac{127}{1344}\frac{m}{\hbar}g^{3}v_{r}T^{8}S_{\text{zzzz}}$ $7.25\times 10^{-10}$ [Uncaptioned image]
15 $\frac{9}{20}\frac{m}{\hbar}gv_{r}v_{z}T^{7}Q_{\text{zzz}}T_{\text{zz}}$ $5.29\times 10^{-10}$ [Uncaptioned image]
16 $\frac{127}{1120}\frac{m}{\hbar}g^{2}v_{r}T^{8}Q_{\text{zzz}}T_{\text{zz}}$ $4.35\times 10^{-10}$ [Uncaptioned image]
17 $\frac{31}{60}\frac{m}{\hbar}gv_{r}v_{z}^{2}T^{6}S_{\text{zzzz}}$ $3.72\times 10^{-10}$ [Uncaptioned image]
18 $\frac{31}{72}\frac{m}{\hbar}v_{r}v_{z}^{2}T^{6}Q_{\text{zzz}}T_{\text{zz}}$ $1.55\times 10^{-10}$ [Uncaptioned image]
19 $\frac{1}{6}\frac{m}{\hbar}v_{r}^{3}T^{4}Q_{\text{zzz}}$ $9.38\times 10^{-11}$ [Uncaptioned image]
20 $\frac{1}{4}\frac{m}{\hbar}v_{r}v_{z}^{3}T^{5}S_{\text{zzzz}}$ $5.50\times 10^{-11}$ [Uncaptioned image]
21 $\frac{1}{40}\frac{m}{\hbar}v_{r}v_{z}T^{7}T_{\text{zz}}^{3}$ $1.96\times 10^{-11}$ [Uncaptioned image]
22 $\frac{127}{20160}\frac{m}{\hbar}gv_{r}T^{8}T_{\text{zz}}^{3}$ $1.61\times 10^{-11}$ [Uncaptioned image]
23 $\frac{1}{4}\frac{1}{\beta_{z}}v_{r}T^{2}Q_{\text{zzz}}$ $6.50\times 10^{-12}$ [Uncaptioned image]
24 $\frac{3}{16}\frac{m}{\hbar}g^{2}v_{r}^{2}T^{7}S_{\text{zzzz}}$ $9.66\times 10^{-13}$ [Uncaptioned image]
25 $\frac{31}{60}\frac{m}{\hbar}gv_{r}^{2}v_{z}T^{6}S_{\text{zzzz}}$ $8.15\times 10^{-13}$ [Uncaptioned image]
26 $\frac{1}{4}\frac{1}{\beta_{x}}v_{r}T^{2}Q_{\text{xxz}}$ $8.13\times 10^{-13}$ [Uncaptioned image]
27 $\frac{1}{4}\frac{1}{\beta_{y}}v_{r}T^{2}Q_{\text{yyz}}$ $8.13\times 10^{-13}$ [Uncaptioned image]
28 $\frac{9}{40}\frac{m}{\hbar}gv_{r}^{2}T^{7}Q_{\text{zzz}}T_{\text{zz}}$ $5.80\times 10^{-13}$ [Uncaptioned image]
29 $\frac{31}{72}\frac{m}{\hbar}v_{r}^{2}v_{z}T^{6}Q_{\text{zzz}}T_{\text{zz}}$ $3.40\times 10^{-13}$ [Uncaptioned image]
30 $\frac{3}{8}\frac{m}{\hbar}v_{r}^{2}v_{z}^{2}T^{5}S_{\text{zzzz}}$ $1.81\times 10^{-13}$ [Uncaptioned image]
@@ ACKNOWLEDGMENTS
We dedicate this paper to Ernst Maria Rasel in celebration of his 60th birthday. His groundbreaking work to broadly push the boundaries of atom interferometry, through advances such as pioneering atom interferometry in microgravity and inventing many new atom optics techniques, is a constant source of inspiration. We thank Cass Sackett and Michael Kagan for valuable discussions. This work was produced by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. The United States Government retains and the publisher, by accepting the work for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this work, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. This work is funded in part by the Gordon and Betty Moore Foundation (Grant GBMF7945), the David and Lucile Packard Foundation (Fellowship for Science and Engineering), and the Office of Naval Research (Grant Number N00014-19-1-2181). JG acknowledges support from a National Science Foundation (NSF) Quantum Information Science and Engineering Network (QISE-NET) Graduate Fellowship, funded by NSF award No. DMR-1747426.

### v3

@@ III.1.1 Diagram Values
where $T_{ij}$, $Q_{ijk}$, and $S_{ijkl}$ are the coefficients of the Taylor series expansion of Earth’s gravitational field in space, evaluated about the the location of the interferometer on the surface of the Earth. We choose the $z$ axis to be the axis of the interferometer. More generally, these coefficients could correspond to the Taylor expansion of any spatially varying potential, such as that generated by a dipole beam, or nonuniform magnetic fields. With $L_{1}=0$, the interferometer phase can be computed by evaluating the 3D version of the spatial integrals of Sec II.2, which results in
@@ III.1.2 Validating with Numerics in 1D
To explore the extent to which these higher order quantum corrections under power law potentials associated with the gravitational field of the Earth could be measured with a state-of-the-art matter wave interferometer, we take the analytic terms that come out of the diagrammatic formalism, and plug in typical operating parameters (listed in the caption of Table 1). The terms are ranked in Table 1 by the size of their contribution to the interferometer phase shift, with the higher-order quantum correction terms highlighted in gray.
The numerical values of the Taylor series coefficients $T_{ij}$, $Q_{ijk}$, and $S_{ijkl}$ ultimately reflect the gravitational potential local to the interferometer and therefore may differ, for a given site and a given time, from the values considered here. The values of the coefficients used in Table 1 are obtained by first assuming a spherical Earth with homogeneous mass density. We then add corrections (also called gravitational anomalies) $T_{a}$, $Q_{a}$, and $S_{a}$ to the coefficients $T_{zz}$, $Q_{zzz}$, and $S_{zzzz}$. Under this prescription, the corrected coefficients become $T_{zz}=-2g/R-T_{a}$, $Q_{zzz}=6g/R^{2}+Q_{a}$, $S_{zzzz}=-24g/R^{3}-S_{a}$. We take the following example values based on the estimates in Ref. [dubetsky_comment_2020]: $T_{a}=10^{-7}~\text{s}^{-2}$, $Q_{a}=3\times 10^{-11}~\text{m}^{-1}\text{s}^{-2}$, $S_{a}=10^{-14}~\text{m}^{-2}\text{s}^{-2}$, $g=9.8~\text{m}/\text{s}^{2}$, and $R=6.38\times 10^{6}~\text{m}$. The coefficients governing the transverse dependence of the gravitational field (e.g. $Q_{xxz}$) are chosen such that the resulting potential satisfies Laplace’s equation. The numerical value of the Taylor series coefficients at a specific interferometer site will differ due to local contributions of the non-spherical nature of the Earth, the inhomogeneous mass density of the Earth, nearby structures, sea level variations, and other site-dependent sources of gravitational potentials [ufrecht_reply_2020]. We emphasize that the calculation presented here is intended as a demonstration of the method for computing quantum corrections to power-law potentials in general. In practice, the true coefficients ($T_{ij}$, $Q_{ijk}$, ...) would need to be measured at the site of the interferometer.
@@
It is important to carry out these calculations in three dimensions. Even though the semiclassical trajectories do not deflect in the directions transverse to the interferometer axis (i.e. the classical paths lie entirely along $z$), quantum corrections still depend on the transverse gravity gradients. For the potentials considered in Table 1, if the initial wavepacket is spherically symmetric ($w_{0,1}=w_{0,2}=w_{0,3}$), then the value of the 1-loop 3-point vertex diagram of Table 1 vanishes: the terms proportional to $Q_{xxz}$ (term 27) and $Q_{yyz}$ (term 28) cancel with the corresponding $Q_{zzz}$ contribution (term 24). This cancellation was noted in Ref. [ufrecht2020perturbative]. The diagrammatic formalism can be used to prove that, more generally, if the wavepacket is prepared with spherical symmetry, then the first-order quantum corrections to the semiclassical approximation vanish for any external potential $V[\mathbf{r}]$ that satisfies Laplace’s equation, $\nabla^{2}V[\mathbf{r}]=0$. We work out this result explicitly in Sec. E. Since the gravitational potential in mass-free regions satisfies Laplace’s equation, it follows that if the initial wavepacket is spherically symmetric, there are no quantum corrections to the interferometer phase shift to first order in $V[\mathbf{r}]$. For the numerical values of Table 1, we take the initial wavepacket to be twice as large along the interferometer axis than in the transverse dimensions, which gives rise to nonzero quantum corrections at first order in $Q_{ijk}$. The asymmetry in the initial wavepacket can be interpreted as arising from asymmetric cooling. For the parameters of Table 1, the momentum-space width is smaller along the interferometer axis than in the transverse directions, corresponding to a lower temperature along that axis. In Ref. [ufrecht2020perturbative], the authors compute the quantum corrections arising from a first-order expansion of the interferometer phase shift in $Q_{ijk}$. Using the same choice of initial wavepacket waists, $\beta_{1}=\beta_{2}=2\beta_{3}=\omega$, our first order result in $Q_{ijk}$ agrees with their first-order expression.
@@
Table 1: Phase shift terms for a Mach-Zehnder matter wave interferometer in 3D subject to the spatially varying gravitational field of the Earth (including gravitational anomalies, see main text for details), along with the diagrams which produce the term, ranked in order of the magnitude of their contribution to the total phase shift. The phase shifts are computed with the following parameters: $k_{z}=2\pi/(698\text{ nm})$, $T=1\text{ s}$, $g=9.8\text{ m/s}^{2}$, $T_{zz}\approx-3.17\times 10^{-6}~\text{s}^{-1}$, $T_{xx}=T_{yy}=-T_{zz}/2$, $Q_{zzz}\approx 3.15\times 10^{-11}~\text{m}^{-1}\text{s}^{-2}$, $Q_{xxz}=Q_{yyz}=-Q_{zzz}/2$, $S_{zzzz}\approx-1.00\times 10^{-14}~\text{m}^{-2}\text{s}^{-2}$, $x_{0}=y_{0}=z_{0}=0$, $v_{z}=3\text{ m/s}$, $v_{x}=v_{y}=0$, $m=86.91\text{u}$, $v_{r}=\hbar k_{z}/m$, $\beta_{\ell}=2\hbar/(mw_{0,\ell}^{2})$ with $\ell=x,y,z$. $w_{0,z}=2\text{ mm}$, $w_{0,x}=w_{0,y}=1\text{ mm}$. The quantum correction terms are highlighted in light gray.
@@
2 $\frac{7}{12}\frac{m}{\hbar}gv_{r}T^{4}T_{\text{zz}}$ $1.63\times 10^{2}$ [Uncaptioned image]
3 $\frac{m}{\hbar}v_{r}v_{z}T^{3}T_{\text{zz}}$ $8.57\times 10^{1}$ [Uncaptioned image]
4 $\frac{1}{2}\frac{m}{\hbar}v_{r}^{2}T^{3}T_{\text{zz}}$ $9.39\times 10^{-2}$ [Uncaptioned image]
5 $\frac{31}{120}\frac{m}{\hbar}g^{2}v_{r}T^{6}Q_{\text{zzz}}$ $7.02\times 10^{-3}$ [Uncaptioned image]
6 $\frac{3}{4}\frac{m}{\hbar}gv_{r}v_{z}T^{5}Q_{\text{zzz}}$ $6.24\times 10^{-3}$ [Uncaptioned image]
7 $\frac{7}{12}\frac{m}{\hbar}v_{r}v_{z}^{2}T^{4}Q_{\text{zzz}}$ $1.49\times 10^{-3}$ [Uncaptioned image]
8 $\frac{31}{360}\frac{m}{\hbar}gv_{r}T^{6}T_{\text{zz}}^{2}$ $7.64\times 10^{-5}$ [Uncaptioned image]
9 $\frac{1}{4}\frac{m}{\hbar}v_{r}v_{z}T^{5}T_{\text{zz}}^{2}$ $6.79\times 10^{-5}$ [Uncaptioned image]
10 $\frac{3}{8}\frac{m}{\hbar}g^{2}v_{r}v_{z}T^{7}S_{\text{zzzz}}$ $9.73\times 10^{-6}$ [Uncaptioned image]
11 $\frac{127}{1344}\frac{m}{\hbar}g^{3}v_{r}T^{8}S_{\text{zzzz}}$ $8.01\times 10^{-6}$ [Uncaptioned image]
12 $\frac{3}{8}\frac{m}{\hbar}gv_{r}^{2}T^{5}Q_{\text{zzz}}$ $6.84\times 10^{-6}$ [Uncaptioned image]
13 $\frac{31}{60}\frac{m}{\hbar}gv_{r}v_{z}^{2}T^{6}S_{\text{zzzz}}$ $4.10\times 10^{-6}$ [Uncaptioned image]
14 $\frac{7}{12}\frac{m}{\hbar}v_{r}^{2}v_{z}T^{4}Q_{\text{zzz}}$ $3.26\times 10^{-6}$ [Uncaptioned image]
15 $\frac{1}{4}\frac{m}{\hbar}v_{r}v_{z}^{3}T^{5}S_{\text{zzzz}}$ $6.08\times 10^{-7}$ [Uncaptioned image]
16 $\frac{1}{8}\frac{m}{\hbar}v_{r}^{2}T^{5}T_{\text{zz}}^{2}$ $7.45\times 10^{-8}$ [Uncaptioned image]
17 $\frac{9}{20}\frac{m}{\hbar}gv_{r}v_{z}T^{7}Q_{\text{zzz}}T_{\text{zz}}$ $1.19\times 10^{-8}$ [Uncaptioned image]
18 $\frac{3}{16}\frac{m}{\hbar}g^{2}v_{r}^{2}T^{7}S_{\text{zzzz}}$ $1.07\times 10^{-8}$ [Uncaptioned image]
19 $\frac{127}{1120}\frac{m}{\hbar}g^{2}v_{r}T^{8}Q_{\text{zzz}}T_{\text{zz}}$ $9.78\times 10^{-9}$ [Uncaptioned image]
20 $\frac{31}{60}\frac{m}{\hbar}gv_{r}^{2}v_{z}T^{6}S_{\text{zzzz}}$ $9.00\times 10^{-9}$ [Uncaptioned image]
21 $\frac{31}{72}\frac{m}{\hbar}v_{r}v_{z}^{2}T^{6}Q_{\text{zzz}}T_{\text{zz}}$ $3.48\times 10^{-9}$ [Uncaptioned image]
22 $\frac{1}{6}\frac{m}{\hbar}v_{r}^{3}T^{4}Q_{\text{zzz}}$ $2.04\times 10^{-9}$ [Uncaptioned image]
23 $\frac{3}{8}\frac{m}{\hbar}v_{r}^{2}v_{z}^{2}T^{5}S_{\text{zzzz}}$ $2.00\times 10^{-9}$ [Uncaptioned image]
24 $\frac{1}{4}\frac{1}{\beta_{z}}v_{r}T^{2}Q_{\text{zzz}}$ $1.42\times 10^{-10}$ [Uncaptioned image]
25 $\frac{1}{40}\frac{m}{\hbar}v_{r}v_{z}T^{7}T_{\text{zz}}^{3}$ $2.15\times 10^{-11}$ [Uncaptioned image]
26 $\frac{127}{20160}\frac{m}{\hbar}gv_{r}T^{8}T_{\text{zz}}^{3}$ $1.77\times 10^{-11}$ [Uncaptioned image]
27 $\frac{1}{4}\frac{1}{\beta_{x}}v_{r}T^{2}Q_{\text{xxz}}$ $1.77\times 10^{-11}$ [Uncaptioned image]
28 $\frac{1}{4}\frac{1}{\beta_{y}}v_{r}T^{2}Q_{\text{yyz}}$ $1.77\times 10^{-11}$ [Uncaptioned image]
29 $\frac{9}{40}\frac{m}{\hbar}gv_{r}^{2}T^{7}Q_{\text{zzz}}T_{\text{zz}}$ $1.30\times 10^{-11}$ [Uncaptioned image]
30 $\frac{31}{72}\frac{m}{\hbar}v_{r}^{2}v_{z}T^{6}Q_{\text{zzz}}T_{\text{zz}}$ $7.63\times 10^{-12}$ [Uncaptioned image]
@@ ACKNOWLEDGMENTS
We dedicate this paper to Ernst Maria Rasel in celebration of his 60th birthday. His groundbreaking work to broadly push the boundaries of atom interferometry, through advances such as pioneering atom interferometry in microgravity and inventing many new atom optics techniques, is a constant source of inspiration. We thank Cass Sackett, Boris Dubetsky, and Michael Kagan for valuable discussions. This work was produced by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. The United States Government retains and the publisher, by accepting the work for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this work, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. This work is funded in part by the Gordon and Betty Moore Foundation (Grant GBMF7945), the David and Lucile Packard Foundation (Fellowship for Science and Engineering), and the Office of Naval Research (Grant Number N00014-19-1-2181). JG acknowledges support from a National Science Foundation (NSF) Quantum Information Science and Engineering Network (QISE-NET) Graduate Fellowship, funded by NSF award No. DMR-1747426.